40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 12 00:14:00 2013, yuta, Update, Locking, PRM coil balancing 
    Reply  Sat Feb 16 17:27:26 2013, yuta, Update, SUS, PRM coil balanced PRM_f2a.png
       Reply  Sun Feb 17 19:27:19 2013, rana, Update, SUS, PRM coil balanced 
       Reply  Wed Feb 27 11:59:43 2013, yuta, Update, SUS, PRM/BS coil re-balanced PRM_f2a20130227.pngBS_f2a20130227.png
Message ID: 8093     Entry time: Sat Feb 16 17:27:26 2013     In reply to: 8065     Reply to this: 8096   8182
Author: yuta 
Type: Update 
Category: SUS 
Subject: PRM coil balanced 

PRM coil gains and f2a filters are adjusted for PRMI work.
It seems like UR/LL coil gains were ~10 % larger than others, and f2a filters changed by few %.

What I did:
  1. Tried to lock PRMI but when I turn on PRCL lock, PRM reflection looked like it tends to go up and left in REFL camera (last night).

  2. So, I set up PRM oplev back, by steering PRM oplev mirrors on the BS table (last night).

  3. Turned PRM oplev sero on, f2a filters off, and ran

> /opt/rtcds/caltech/c1/scripts/SUS/F2P_LOCKIN.py -o PRM

  I had to fix F2P_LOCKIN.py because it assumed some OUTPUT buttons in LOCKIN1 filters to be ON.
  Also, I had to restore filters in LOCKIN1 (8.5 Hz bandpass filter etc.) because their names were changed. To do this, I copied filters needed from /opt/rtcds/caltech/c1/chans/filter_archive/c1sus/C1SUS_110916_162512.txt, renamed LOCKIN1_(I|Q|SIG) with LOCKIN1_DEMOD_(I|Q|SIG), and pasted to the current filter bank file. I checked that they look OK with foton after editing the file.

  This measurement takes about 30 minutes. I ran several times to check consistency. There was ~ 0.1 % standard deviation for the measurement results.

  4. By putting measured coupling coefficients and PRM pendulum frequency (f0=0.993 Hz) to /opt/rtcds/caltech/c1/scripts/SUS/F2Pcalc.py, I got new f2a filters.

  5. Overwrote f2a filters in C1:SUS-PRM_TO_COIL_(1-4)_1 FM1 with new ones, and turned  new f2a filters on.

Result:
  Below is the DC gain adjustment result from F2P_LOCKIN.py;

multiplier factors are :
UL = 1.141525
UR = 0.879997
LR = 1.117484
LL = 0.860995
Set C1:SUS-PRM_ULCOIL_GAIN to 1.04990177238
Set C1:SUS-PRM_URCOIL_GAIN to -0.983396190716
Set C1:SUS-PRM_LRCOIL_GAIN to 0.954304254663
Set C1:SUS-PRM_LLCOIL_GAIN to -0.971356852259


  So, UR/LL coil gains somehow got ~10 % larger than other two since last coil balancing.

  Measured coupling coefficients from F2P_LOCKIN.py were

- measured coupling coefficients are :
P2P(POS=>PIT) = 0.014993
P2Y(POS=>YAW) = 0.001363


  New f2a filters are plotted below. They look fairly different compared with previous ones.
PRM_f2a.png


 

We need better F2P_LOCKIN.py:
  Some one should make F2P_LOCKIN.py better. The main problem is the sudden gain change when starting diagonalization at low frequency. It sometimes trips off the watchdog.

Some elogs related:
  Kiwamu made f2a filters in Sep 2011: elog #5417
  Koji adjusting DC gains in Jan 2013: elog #7969

ELOG V3.1.3-